kisha has five different marbles. How many different ways can she place the marbles into two bags so that at least one marble is in each bag
a 15
b 60
c 30
d 25

Answer by stanbon(57219)

kisha has five different marbles. How many d
ifferent ways can she place the marbles into two bags so that at least one marble is in each bag
---
pick a bag: 2 ways
Put at least 1 marble in the bag:
Note: each of the 5 marbles is either in or not in the 1st bag: 2^5 = 32 ways
But the bag cannot have 0 marbles and cannot have all 5 marbles.
2^5-2 = 30 ways
-----
Ans: 2*30 = 60 ways
=========================
Cheers,
Stan H.
=========================
a 15
b 60
c 30
d 25